OpenStudy (anonymous):

Simplify. -8 + (-6)x - 5 + 2x

OpenStudy (anonymous):

= -8 + (-6)x - 5 +2x = -8 - 6x - 5 + 2x = -4x - 13

OpenStudy (anonymous):

-13-4x